# The KSBA moduli space of stable log Calabi–Yau surfaces

### 23 Feb 2024 v1

@@ 4.2. Open Kulikov surfaces from polarized log Calabi-Yau surfaces
STEP I: As a first step, for every vertex $v\in\mathscr{P}^{[0]}$, we define a toric surface $\widetilde{X}^{v}_{\mathscr{P}}$ by deformation of the union $\bigcup_{v^{\prime}\in pr^{-1}(v)}X^{v}_{\overline{\mathscr{P}}}$ of irreducible components of $\mathcal{X}_{\overline{\mathscr{P}},0}$, where $pr:\overline{\mathscr{P}}^{[0]}\rightarrow\mathscr{P}^{[0]}$ is the natural projection coming from the description of the Symington of $P$ as a quotient of $\overline{P}\setminus\cup_{ij}\mathrm{Int}(\Delta_{ij})$. If $pr^{-1}(v)$ is of cardinality one, with $pr^{-1}(v)=\{v^{\prime}\}$, then we set simply $\widetilde{X}^{v}_{\mathscr{P}}:=X^{v^{\prime}}_{\overline{\mathscr{P}}}$.
If $pr^{-1}(v)$ is of cardinality strictly bigger than one, then $pr^{-1}(v)=\{v^{\prime}_{1},\dots,v^{\prime\prime}_{r}\}$, where for every $11\leq k\leq r-$ 1, $v^{\prime}_{k}$ and $v^{\prime}_{k+1}$ are opposite integral points on the interior edges of a triangle $\Delta_{i_{k}j_{k}}$. As the good decomposition $\overline{\mathscr{P}}$ is a standard decomposition in restriction to $\Delta_{ij}$, as in Definition 2.13, it follows that each toric surface $X^{v_{k}^{\prime}}_{\overline{\mathscr{P}}}$ has a map to $\mathbb{P}^{1}$ induced by the opposite rays in its fan coming from the interior edge of $\Delta_{i_{k}j_{k}}$ or $\Delta_{i_{k-1}j_{k-1}}$ containing $v_{k}^{\prime}$. These maps $f_{k}:X^{v_{k}^{\prime}}_{\overline{\mathscr{P}}}\rightarrow\mathbb{P}^{1}$ are generically $\mathbb{P}^{1}$-fibrations. Moreover, for every $11\leq k\leq r-$ 1, there is an edge in $\mathscr{P}^{[1]}$ contained in $\Delta_{i_{k}j_{k}}$ and connecting the vertices $v_{k}^{\prime}$ and $v_{k+1}^{\prime}$. This edge corresponds to the intersection $F_{k}:=X^{v_{k}^{\prime}}_{\overline{\mathscr{P}}}\cap X^{v_{k+1}^{\prime}}_{\overline{\mathscr{P}}}$, and is both a $\mathbb{P}^{1}$-fiber of $f_{k}$ in $X^{v_{k}^{\prime}}_{\overline{\mathscr{P}}}$ and a $\mathbb{P}^{1}$-fiber of $f_{k+1}$ in $X^{v_{k+1}^{\prime}}_{\overline{\mathscr{P}}}$.
@@
Hence, $\bigcup_{i=1}^{k}X^{v_{k}^{\prime}}_{\overline{\mathscr{P}}}$ is a chain of generically $\mathbb{P}^{1}$-fibrations, glued along $\mathbb{P}^{1}$-fibers. The double locus of such a normal crossing union of toric surfaces can be smoothed by a toric degeneration defined by a polyhedral decomposition of $\mathbb{R}^{2}$ as in Figure 4.1. We define $\widetilde{\mathcal{X}}_{\mathscr{P}}^{v}$ as the toric surface given by the general fiber of this deformation. The fan of $\widetilde{\mathcal{X}}_{\mathscr{P}}^{v}$ can be described explicitly as follows. First we draw the fan of $X^{v_{1}^{\prime}}_{\overline{\mathscr{P}}}$ such that the opposite rays defining $f_{1}$ are horizontal and such that the ray corresponding to $F_{1}$ is contained in the upper-half plane. We also draw the fan of $X^{v_{r}^{\prime}}_{\overline{\mathscr{P}}}$ such that the opposite rays defining $f_{r}$ are horizontal and such that the ray corresponding to $F_{r-1}$ is contained in the lower-half plane. Then, the fan of $\widetilde{\mathcal{X}}_{\mathscr{P}}^{v}$ is equal to the fan of $X^{v_{1}^{\prime}}_{\overline{\mathscr{P}}}$ in restriction to the lower half-plane, and equal to the fan of $X^{v_{r}^{\prime}}_{\overline{\mathscr{P}}}$ in restriction to the upper half-plane.
STEP II: As a second step, for each vertex $v\in\mathscr{P}^{[0]}$, we describe a deformation of the toric pair $(\widetilde{X}_{\mathscr{P}}^{v},\partial\widetilde{X}_{\overline{\mathscr{P}}}^{v})$ into a log Calabi–Yau pair $(X_{\overline{\mathscr{P}}}^{v},\partial X_{\overline{\mathscr{P}}}^{v})$. Recall by Construction 2.10, the Symington polytope $P$ is obtained from $\overline{P}$, by cutting out triangles $\Delta_{ij}$, with one edge in the boundary of $\overline{P}$ and with an interior vertex $p_{ij}$. For every $i,j$ such that $v=pr(p_{ij})$, there is a face $F_{ij}$ of $\overline{\mathscr{P}}$ containing $v$ and contained in the triangle $\Delta_{ij}$. The face $F_{ij}$ defines a 2-dimensional cone in the star of $\overline{\mathscr{P}}$ at $v$, that is, in the fan of $X_{\overline{\mathscr{P}}}^{v}$, and so corresponds to a toric fixed point $x_{ij}$ of $X_{\overline{\mathscr{P}}}^{v}$. This point is not smoothed by the deformation in Step I, and so also defines a toric fixed point in the toric surface $\widetilde{X}_{\mathscr{P}}^{v}$ that we still denote by $x_{ij}$. We denote by $(V_{ij},\partial V_{ij})$ the analytic germ of the surface pair $(\widetilde{X}_{\mathscr{P}}^{v},\partial\widetilde{X}_{\mathscr{P}}^{v})$ around $x_{ij}$. As the triangles $\Delta_{ij}$ have equal lattice height and lattice base, it follows from [11, §7.3] that $(V_{ij},\partial V_{ij})$ is the germ of the origin in an affine toric variety of the form
@@
Let $x$ be a 00-dimensional stratum of $D$. Then, $\overline{X}^{v_{x}}_{\mathscr{P}}$ is an affine surface obtained by a flat deformation of the affine toric surface $\mathrm{Spec}\,\mathbb{C}[\sigma_{x}\cap\mathbb{Z}^{2}]$, and $\partial\overline{X}^{v_{x}}_{\mathscr{P}}$ is a flat locally trivial deformation of the union of the two toric divisors of $\mathrm{Spec}\,\mathbb{C}[\sigma_{x}\cap\mathbb{Z}^{2}]$. Moreover, the affinization map $X^{v_{x}}_{\mathscr{P}}\rightarrow\overline{X}^{v_{x}}_{\mathscr{P}}$ is a surjective projective birational morphism.
@@
In general, if there are points $p_{ij}$ such that $pr(p_{ij})=v_{x}$, then, by Step II of the construction of $\mathcal{X}_{\mathscr{P},0}$ in §4.2, $X^{v_{x}}_{\mathscr{P}}$ is a flat deformation, obtained by partial smoothing of the toric boundary, of the toric surface $X^{v_{x}}_{\overline{\mathscr{P}}}$ whose fan is given by the star of $\mathscr{P}$ at $v_{x}$, and whose affinization is given by $\overline{X}^{v_{x}}_{\overline{\mathscr{P}}}=\mathrm{Spec}\,\mathbb{C}[\sigma_{x}\cap\mathbb{Z}^{2}]$. As $X^{v_{x}}_{\overline{\mathscr{P}}}$ is a toric surface, we have $H^{1}(X^{v_{x}}_{\overline{\mathscr{P}}},\mathcal{O}_{X^{v_{x}}_{\overline{\mathscr{P}}}})=0$. Hence, by [95, Theorem 1.4] – see also [26, Theorem 3.1], the affinization $X^{v_{x}}_{\mathscr{P}}\rightarrow\overline{X}^{v_{x}}_{\mathscr{P}}$ is a flat deformation of the toric affinization $X^{v_{x}}_{\overline{\mathscr{P}}}\rightarrow\overline{X}^{v_{x}}_{\overline{\mathscr{P}}}=\mathrm{Spec}\,\mathbb{C}[\sigma_{x}\cap\mathbb{Z}^{2}]$, and this implies Lemma 4.3. ∎
@@
Let $v\in(\partial P)_{\mathbb{Z}}$ such that $v\neq v_{x}$ for all 00-dimensional strata $x$ of $D$. Then, we have $\overline{X}^{v}_{\mathscr{P}}\simeq\mathbb{A}^{1}$, and the affinization map $X^{v}_{\mathscr{P}}\rightarrow\overline{X}^{v}_{\mathscr{P}}\simeq\mathbb{A}^{1}$ is a surjective projective morphism with general fiber $\mathbb{P}^{1}$.
@@
In general, if there are points $p_{ij}$ such that $pr(p_{ij})=v$, then, by Step II of the construction of $\mathcal{X}_{\mathscr{P},0}$ in §4.2, $X^{v}_{\mathscr{P}}$ is a flat deformation, obtained by partial smoothing of the toric boundary, of the toric surface $X^{v}_{\overline{\mathscr{P}}}$ whose fan is given by the star of $\mathscr{P}$ at $v$. As in the previous case, we have a surjective projective toric morphism $X^{v}_{\overline{\mathscr{P}}}\rightarrow\mathbb{A}^{1}$ with general fiber $\mathbb{P}^{1}$. Since $X^{v}_{\overline{\mathscr{P}}}$ is a toric surface, we have $H^{1}(X^{v}_{\overline{\mathscr{P}}},\mathcal{O}_{X^{v}_{\overline{\mathscr{P}}}})=0$, and so by [95, Theorem 1.4], the morphism $X^{v}_{\overline{\mathscr{P}}}\rightarrow\mathbb{A}^{1}$ deforms to a surjective projective morphism $X^{v}_{\mathscr{P}}\rightarrow\mathbb{A}^{1}$ with general fiber $\mathbb{P}^{1}$. This morphism is necessarily the affinization map of $X^{v}_{\mathscr{P}}$, and so Lemma 4.3 follows. ∎
@@
When $Y$ is smooth and the central point $p\in\overline{P}$ is in the interior of $\overline{P}$, then we have $\overline{X}^{v_{x}}_{\mathscr{P}}\simeq\mathbb{A}^{2}$ for every 00-dimensional stratum $x$ of $D$, and so $\overline{\mathcal{X}}_{\mathscr{P},0}$ is isomorphic to the $n$-vertex as in [40, §1.3], where $n$ is the number of irreducible components of $D$. In general, by Lemma 4.3, the singularities of the irreducible components $\overline{X}^{v_{x}}_{\mathscr{P}}$ at the central point $x_{0}$ are flat deformations of cyclic quotient singularities, and so are also cyclic quotient singularities by [65, Corollary 7.15].
@@
The normal crossing surface $\mathcal{X}_{\mathscr{P},0}$ is a quasi-projective open Kulikov surface.
@@
It follows from the description of the irreducible components of $\mathcal{X}_{\mathscr{P},0}$ in Lemmas 4.3 and 4.5 that $\mathcal{X}_{\mathscr{P},0}$ is an open Kulikov surface. To prove quasi-projectivity, first recall that, by Definition 2.13, the good polyhedral decomposition $\overline{\mathscr{P}}$ is regular. Hence, there exists a convex piecewise-linear function $\varphi$ on $\overline{P}$ inducing $\overline{\mathscr{P}}$. We denote by $L_{\varphi}$ the corresponding ample line bundle on $\mathcal{X}_{\overline{\mathscr{P}}}$. If follows from the toric description of the deformation in Step I of the construction of $\mathcal{X}_{\mathscr{P},0}$ in §4.2 that the restrictions of $L_{\varphi}$ to $\bigcup_{v^{\prime}\in pr^{-1}(v)}X^{v}_{\overline{\mathscr{P}}}$ deform into ample line bundles $\widetilde{L}_{\varphi}^{v}$ on the toric surfaces $\widetilde{X}^{v}_{\mathscr{P}}$. As a toric surface $X$ satisfies $H^{2}(X,\mathcal{O}_{X})=0$, it follows from deformation theory for line bundles (see for example the proof of [84, Theorem 2.5.23 (ii)]) that, in Step II of the construction of $\mathcal{X}_{\mathscr{P},0}$ in §4.2, the ample line bundles $L^{v}_{\varphi}$ on the toric surfaces $\widetilde{X}^{v}_{\mathscr{P}}$ deform into ample line bundles $L^{v}_{\varphi}$ on the log Calabi–Yau surfaces $X^{v}_{\mathscr{P}}$. Moreover, the line bundles $L^{v}_{\varphi}$ and $L^{v^{\prime}}_{\varphi}$ are isomorphic in restriction to $X^{e}_{\mathscr{P}}$, for every edge $e\in\mathscr{P}^{[1]}$ adjacent to two vertices $v,v^{\prime}\in\mathscr{P}^{[1]}$. Hence, it follows from the description of $\mathrm{Pic}(\mathcal{X}_{\mathscr{P},0})$ in Lemma 3.10 that the ample line bundles $L^{v}_{\varphi}$ glue into a global ample line bundle on $\mathcal{X}_{\mathscr{P},0}$. ∎

### 30 Mar 2024 v2

@@ 4.2. Open Kulikov surfaces from polarized log Calabi-Yau surfaces
STEP I: As a first step, for every vertex $v\in\mathscr{P}^{[0]}$, we define a toric surface $\widetilde{X}^{v}_{\mathscr{P}}$ by deformation of the union $\bigcup_{v^{\prime}\in pr^{-1}(v)}X^{v^{\prime}}_{\overline{\mathscr{P}}}$ of irreducible components of $\mathcal{X}_{\overline{\mathscr{P}},0}$, where $pr:\overline{\mathscr{P}}^{[0]}\rightarrow\mathscr{P}^{[0]}$ is the natural projection coming from the description of the Symington of $P$ as a quotient of $\overline{P}\setminus\cup_{ij}\mathrm{Int}(\Delta_{ij})$. If $pr^{-1}(v)$ is of cardinality one, with $pr^{-1}(v)=\{v^{\prime}\}$, then we set simply $\widetilde{X}^{v}_{\mathscr{P}}:=X^{v^{\prime}}_{\overline{\mathscr{P}}}$. If $pr^{-1}(v)$ is of cardinality strictly bigger than one, then $pr^{-1}(v)=\{v^{\prime}_{1},\dots,v^{\prime}_{r}\}$, where for every $11\leq k\leq r-$ 1, $v^{\prime}_{k}$ and $v^{\prime}_{k+1}$ are opposite integral points on the interior edges of a triangle $\Delta_{i_{k}j_{k}}$. As the good decomposition $\overline{\mathscr{P}}$ is a standard decomposition in restriction to $\Delta_{ij}$, as in Definition 2.13, it follows that each toric surface $X^{v_{k}^{\prime}}_{\overline{\mathscr{P}}}$ has a map to $\mathbb{P}^{1}$ induced by the opposite rays in its fan coming from the interior edge of $\Delta_{i_{k}j_{k}}$ or $\Delta_{i_{k-1}j_{k-1}}$ containing $v_{k}^{\prime}$. These maps $f_{k}:X^{v_{k}^{\prime}}_{\overline{\mathscr{P}}}\rightarrow\mathbb{P}^{1}$ are generically $\mathbb{P}^{1}$-fibrations if $v_{k}^{\prime}$ is contained in the interior of $\overline{P}$, and generically $\mathbb{A}^{1}$-fibrations if $v_{k}^{\prime}$ is on the boundary of $\overline{P}$. Moreover, for every $11\leq k\leq r-$ 1, there is an edge in $\mathscr{P}^{[1]}$ contained in $\Delta_{i_{k}j_{k}}$ and connecting the vertices $v_{k}^{\prime}$ and $v_{k+1}^{\prime}$. This edge corresponds to the intersection $F_{k}:=X^{v_{k}^{\prime}}_{\overline{\mathscr{P}}}\cap X^{v_{k+1}^{\prime}}_{\overline{\mathscr{P}}}$, and is both a fiber of $f_{k}$ in $X^{v_{k}^{\prime}}_{\overline{\mathscr{P}}}$ and a fiber of $f_{k+1}$ in $X^{v_{k+1}^{\prime}}_{\overline{\mathscr{P}}}$.
@@
Hence, $\bigcup_{i=1}^{k}X^{v_{k}^{\prime}}_{\overline{\mathscr{P}}}$ is a chain of fibered surfaces, glued along common fibers. The double locus of such a normal crossing union of toric surfaces can be smoothed by a toric degeneration defined by a polyhedral decomposition of $\mathbb{R}^{2}$ as in Figure 4.1. We define $\widetilde{\mathcal{X}}_{\mathscr{P}}^{v}$ as the toric surface given by the general fiber of this deformation. The fan of $\widetilde{\mathcal{X}}_{\mathscr{P}}^{v}$ can be described explicitly as follows. First we draw the fan of $X^{v_{1}^{\prime}}_{\overline{\mathscr{P}}}$ such that the opposite rays defining $f_{1}$ are horizontal and such that the ray corresponding to $F_{1}$ is contained in the upper-half plane. We also draw the fan of $X^{v_{r}^{\prime}}_{\overline{\mathscr{P}}}$ such that the opposite rays defining $f_{r}$ are horizontal and such that the ray corresponding to $F_{r-1}$ is contained in the lower-half plane and parallel to the ray corresponding to $F_{1}$ in the fan of $X^{v_{1}^{\prime}}_{\overline{\mathscr{P}}}$. Then, the fan of $\widetilde{\mathcal{X}}_{\mathscr{P}}^{v}$ is equal to the fan of $X^{v_{1}^{\prime}}_{\overline{\mathscr{P}}}$ in restriction to the lower half-plane, and equal to the fan of $X^{v_{r}^{\prime}}_{\overline{\mathscr{P}}}$ in restriction to the upper half-plane.
STEP II: As a second step, for each vertex $v\in\mathscr{P}^{[0]}$, we describe a deformation of the toric pair $(\widetilde{X}_{\mathscr{P}}^{v},\partial\widetilde{X}_{\mathscr{P}}^{v})$ into a log Calabi–Yau pair $(X_{\mathscr{P}}^{v},\partial X_{\mathscr{P}}^{v})$. Recall by Construction 2.10, the Symington polytope $P$ is obtained from $\overline{P}$, by cutting out triangles $\Delta_{ij}$, with one edge in the boundary of $\overline{P}$ and with an interior vertex $p_{ij}$. For every $i,j$ such that $v=pr(p_{ij})$, there is a face $F_{ij}$ of $\overline{\mathscr{P}}$ containing $v$ and contained in the triangle $\Delta_{ij}$. The face $F_{ij}$ defines a 2-dimensional cone in the star of $\overline{\mathscr{P}}$ at $v$, that is, in the fan of $X_{\overline{\mathscr{P}}}^{v}$, and so corresponds to a toric fixed point $x_{ij}$ of $X_{\overline{\mathscr{P}}}^{v}$. This point is not smoothed by the deformation in Step I, and so also defines a toric fixed point in the toric surface $\widetilde{X}_{\mathscr{P}}^{v}$ that we still denote by $x_{ij}$. We denote by $(V_{ij},\partial V_{ij})$ the analytic germ of the surface pair $(\widetilde{X}_{\mathscr{P}}^{v},\partial\widetilde{X}_{\mathscr{P}}^{v})$ around $x_{ij}$. As the triangles $\Delta_{ij}$ have equal lattice height and lattice base, it follows from [11, §7.3] that $(V_{ij},\partial V_{ij})$ is the germ of the origin in an affine toric variety of the form
@@
Let $(Y,D,L)$ be a polarized log Calabi–Yau surface, with a polyhedral decomposition $\mathscr{P}$ of a Symington polytope $P$ defined as above. For every 00-dimensional stratum of $D$, $\overline{X}^{v_{x}}_{\mathscr{P}}$ is an affine surface obtained by a flat deformation of the affine toric surface $\mathrm{Spec}\,\mathbb{C}[\sigma_{x}\cap\mathbb{Z}^{2}]$, and $\partial\overline{X}^{v_{x}}_{\mathscr{P}}$ is a flat locally trivial deformation of the union of the two toric divisors of $\mathrm{Spec}\,\mathbb{C}[\sigma_{x}\cap\mathbb{Z}^{2}]$. Moreover, the affinization map $X^{v_{x}}_{\mathscr{P}}\rightarrow\overline{X}^{v_{x}}_{\mathscr{P}}$ is a surjective projective birational morphism.
@@
In general, if there are points $p_{ij}$ such that $pr(p_{ij})=v_{x}$, then, by Step II of the construction of $\mathcal{X}_{\mathscr{P},0}$ in §4.2, $X^{v_{x}}_{\mathscr{P}}$ is a flat deformation, obtained by partial smoothing of the toric boundary, of the toric surface $X^{v_{x}}_{\overline{\mathscr{P}}}$ whose fan is given by the star of $\mathscr{P}$ at $v_{x}$, and whose affinization is given by $\overline{X}^{v_{x}}_{\overline{\mathscr{P}}}=\mathrm{Spec}\,\mathbb{C}[\sigma_{x}\cap\mathbb{Z}^{2}]$. As $X^{v_{x}}_{\overline{\mathscr{P}}}$ is a toric surface, we have $H^{1}(X^{v_{x}}_{\overline{\mathscr{P}}},\mathcal{O}_{X^{v_{x}}_{\overline{\mathscr{P}}}})=0$. Hence, by [96, Theorem 1.4] – see also [26, Theorem 3.1], the affinization $X^{v_{x}}_{\mathscr{P}}\rightarrow\overline{X}^{v_{x}}_{\mathscr{P}}$ is a flat deformation of the toric affinization $X^{v_{x}}_{\overline{\mathscr{P}}}\rightarrow\overline{X}^{v_{x}}_{\overline{\mathscr{P}}}=\mathrm{Spec}\,\mathbb{C}[\sigma_{x}\cap\mathbb{Z}^{2}]$, and this implies Lemma 4.3. ∎
@@
Let $(Y,D,L)$ be a polarized log Calabi–Yau surface, with a polyhedral decomposition $\mathscr{P}$ of a Symington polytope $P$ defined as above. For every vertex $v\in(\partial P)_{\mathbb{Z}}$ such that $v\neq v_{x}$ for all 00-dimensional strata $x$ of $D$, we have $\overline{X}^{v}_{\mathscr{P}}\simeq\mathbb{A}^{1}$, and the affinization map $X^{v}_{\mathscr{P}}\rightarrow\overline{X}^{v}_{\mathscr{P}}\simeq\mathbb{A}^{1}$ is a surjective projective morphism with general fiber $\mathbb{P}^{1}$.
@@
In general, if there are points $p_{ij}$ such that $pr(p_{ij})=v$, then, by Step II of the construction of $\mathcal{X}_{\mathscr{P},0}$ in §4.2, $X^{v}_{\mathscr{P}}$ is a flat deformation, obtained by partial smoothing of the toric boundary, of the toric surface $X^{v}_{\overline{\mathscr{P}}}$ whose fan is given by the star of $\mathscr{P}$ at $v$. As in the previous case, we have a surjective projective toric morphism $X^{v}_{\overline{\mathscr{P}}}\rightarrow\mathbb{A}^{1}$ with general fiber $\mathbb{P}^{1}$. Since $X^{v}_{\overline{\mathscr{P}}}$ is a toric surface, we have $H^{1}(X^{v}_{\overline{\mathscr{P}}},\mathcal{O}_{X^{v}_{\overline{\mathscr{P}}}})=0$, and so by [96, Theorem 1.4], the morphism $X^{v}_{\overline{\mathscr{P}}}\rightarrow\mathbb{A}^{1}$ deforms to a surjective projective morphism $X^{v}_{\mathscr{P}}\rightarrow\mathbb{A}^{1}$ with general fiber $\mathbb{P}^{1}$. This morphism is necessarily the affinization map of $X^{v}_{\mathscr{P}}$, and so Lemma 4.3 follows. ∎
@@
When $Y$ is smooth and the central point $p\in\overline{P}$ is in the interior of $\overline{P}$, then we have $\overline{X}^{v_{x}}_{\mathscr{P}}\simeq\mathbb{A}^{2}$ for every 00-dimensional stratum $x$ of $D$, and so $\overline{\mathcal{X}}_{\mathscr{P},0}$ is isomorphic to the $n$-vertex as in [41, §1.3], where $n$ is the number of irreducible components of $D$. In general, by Lemma 4.3, the singularities of the irreducible components $\overline{X}^{v_{x}}_{\mathscr{P}}$ at the central point $x_{0}$ are flat deformations of cyclic quotient singularities, and so are also cyclic quotient singularities by [68, Corollary 7.15].
@@
Let $(Y,D,L)$ be a polarized log Calabi–Yau surface, with a polyhedral decomposition $\mathscr{P}$ of a Symington polytope $P$ defined as above. Then, the normal crossing surface $\mathcal{X}_{\mathscr{P},0}$ is a quasi-projective open Kulikov surface.
@@
It follows from the description of the irreducible components of $\mathcal{X}_{\mathscr{P},0}$ in Lemmas 4.3 and 4.5 that $\mathcal{X}_{\mathscr{P},0}$ is an open Kulikov surface. To prove quasi-projectivity, first recall that, by Definition 2.13, the good polyhedral decomposition $\overline{\mathscr{P}}$ is regular. Hence, there exists a convex piecewise-linear function $\varphi$ on $\overline{P}$ inducing $\overline{\mathscr{P}}$. We denote by $L_{\varphi}$ the corresponding ample line bundle on $\mathcal{X}_{\overline{\mathscr{P}}}$. It follows from the toric description of the deformation in Step I of the construction of $\mathcal{X}_{\mathscr{P},0}$ in §4.2 that the restrictions of $L_{\varphi}$ to $\bigcup_{v^{\prime}\in pr^{-1}(v)}X^{v}_{\overline{\mathscr{P}}}$ deform into ample line bundles $\widetilde{L}_{\varphi}^{v}$ on the toric surfaces $\widetilde{X}^{v}_{\mathscr{P}}$. As a toric surface $X$ satisfies $H^{2}(X,\mathcal{O}_{X})=0$, it follows from deformation theory for line bundles (see for example the proof of [85, Theorem 2.5.23 (ii)]) that, in Step II of the construction of $\mathcal{X}_{\mathscr{P},0}$ in §4.2, the ample line bundles $L^{v}_{\varphi}$ on the toric surfaces $\widetilde{X}^{v}_{\mathscr{P}}$ deform into ample line bundles $L^{v}_{\varphi}$ on the log Calabi–Yau surfaces $X^{v}_{\mathscr{P}}$. Moreover, the line bundles $L^{v}_{\varphi}$ and $L^{v^{\prime}}_{\varphi}$ are isomorphic in restriction to $X^{e}_{\mathscr{P}}$, for every edge $e\in\mathscr{P}^{[1]}$ adjacent to two vertices $v,v^{\prime}\in\mathscr{P}^{[1]}$. Hence, it follows from the description of $\mathrm{Pic}(\mathcal{X}_{\mathscr{P},0})$ in Lemma 3.10 that the ample line bundles $L^{v}_{\varphi}$ glue into a global ample line bundle on $\mathcal{X}_{\mathscr{P},0}$. ∎